how many ways can the teacher select 5 multiple choice and 3 free responses questions from a question bank that contains 15 multiple choice and 6 free responses
Answer by stanbon(75887)

how many ways can the teacher select 5 multiple choice and 3 free responses questions from a question bank that contains 15 multiple choice and 6 free responses
---------
Ans: 15C5 * 6C3 = 3003*20 = 60060
====================================
Cheers,
Stan H.
===================